# A Global Minimum Tax for Large Firms Only:

### v5

@@
Abstract
@@
The Global Minimum Tax (GMT) is applied only to firms above a certain size threshold, permitting countries to set differential tax rates for small and large firms. We analyze tax competition among multiple tax havens and a non-haven country for heterogeneous multinationals to evaluate the effects of this partial coverage of GMT. Upon the introduction of a moderately low GMT rate, the havens commit to the single uniform GMT rate for all multinationals. However, gradual increases in the GMT rate induce the havens, and subsequently the non-haven, to adopt discriminatory, lower tax rates for small multinationals. Our calibration exercise shows that the implementation of a 15% GMT rate results in a regime where only the havens adopt split tax rates. Upon GMT introduction, welfare and tax revenues fall in the tax havens but rise in the non-haven, yielding a positive net gain worldwide.
Keywords: multinational firms; tax avoidance; global minimum tax; profit shifting
JEL classification: F23, H25, H87
1 Introduction
Profit shifting by multinational enterprises (MNEs) has long been a major problem for corporate taxation worldwide. Using aggregate data, Tørsløv et al., (2023) estimate that more than one-third of all foreign-earned corporate profits of multinationals are shifted to tax havens. The important role of tax havens for profit shifting is confirmed in other studies using microdata of multinational affiliates/headquarters in France (Davies et al.,, 2018), the UK (Bilicka,, 2019), and the US (Guvenen et al.,, 2022). In response to the large revenue losses caused by profit shifting, the OECD has launched an action plan to fight base erosion in OECD countries, and in particular the profit shifting to tax havens (OECD,, 2013). A core development in this endeavor is the introduction of a global minimum tax (GMT), which is applied to large MNEs (OECD, 2020a, ; OECD, 2020b, ). The GMT with a tax rate of 15% has been agreed upon by a group of more than 130 countries in 2021 including many tax havens, and most of these countries are expected to have the GMT enacted by 2026.111The list of countries joining the agreement on the GMT, also known as Pillar 2, is found in OECD, (2021). The United States has already enacted, in its 2017 Tax Cuts and Jobs Act, a tax on Global Intangible Low-Tax Income (GILTI) that works in many respects like a GMT (see Chodorow-Reich et al.,, 2024 for details). The global revenue gains from a 15% GMT are estimated to be in the range of 155 to 192 billion USD, or 6.5 to 8.1% of global corporate tax revenues (Hugger et al.,, 2024, Section 8.2). [^2] Baraké et al., (2022, Table 1) break down the global revenue changes of the GMT by country and arrive at a revenue gain of around 55 to 67 billion EUR for the US, and a gain of similar magnitude for the EU as a whole. Estimates of both Hugger et al., (2024) and Baraké et al., (2022) are based on models where governments and multinational firms mechanically respond to the GMT. Wang, (2020) sets up a quantitative corporate tax framework that endogenizes the behavior of both firms and governments. Using similar frameworks, Ferrari et al., (2023) and Shen, (2024) quantify the effects of the GMT. While no econometric estimates exist for the global minimum tax yet, there are experiences from similar policies enacted at the national level. See Buettner and Poehnlein, (2024) for an analysis of the effects of a minimum tax rate for German municipalities.
@@ 1 Introduction
An important limitation of the GMT is, however, that it applies only to “large” multinationals, defined as multinational groups that have had no less than 750 million EUR in total annual revenues in at least two out of the last four years. [^3] Since 2016, multinational groups above this threshold size are also required to file individual reports on their activities, profits, and taxes paid in each of the countries in which they have a presence (country-by-country reporting). Fig. 1 shows that, according to the Orbis database of Bureau van Dijk, about 30% of all MNEs, which are responsible for about 90% of all MNE profits, have consolidated accounts above this threshold and are therefore covered by the GMT. [^4] This corresponds to the estimate of the OECD, which arrives at a coverage rate of 90% of profits, using the Orbis database and other data sources (OECD, 2020a, , # 505, p. 233). Nevertheless, a substantial amount of MNE profits exceeding 300 billion EUR remain outside the scope of the GMT. Moreover, the MNE sample in Orbis is not comprehensive and the database is known for oversampling large firms (Bajgar et al.,, 2020). Therefore, a coverage rate of the GMT of 90% of all MNE profits can be considered as an upper bound of the true profit share covered by the GMT.
One implication of this incomplete coverage is that low-tax countries might respond to the introduction of the GMT by using a split corporate tax system, where a lower tax rate applies to firms below the GMT threshold. In fact, such a split is inherent in the regulation that tax havens are allowed to implement the GMT by raising their (low) general tax rate only for MNEs above the GMT threshold through a specific top-up tax. [^5] This is known as the Qualified Domestic Marginal Top-Up Tax (QDMTT). See Devereux, (2023) for a discussion. Tax havens are fully aware of this option. Ireland and Liechtenstein, for example, have already decided to keep their general tax rates at 12.5%, but top up the tax rate to 15% for affiliates of foreign MNEs above the threshold (Government of Ireland,, 2023; PwC,, 2024 for Liechtenstein). Hungary and Bulgaria have adopted a similarly split tax rate by joining the GMT agreement while maintaining their regular tax rates of 9% and 10%, respectively, for firms not covered by the GMT (Ernst & Young,, 2023 for Hungary; PwC,, 2023 for Bulgaria). [^6] The split corporate tax system is also expected to be introduced in Bermuda and Jersey (Government of Bermuda,, 2024; Government of Jersey,, 2023). These responses indicate that the effects of the GMT cannot be analyzed assuming a single, uniform corporate tax rate across all countries.
Figure 1: GMT coverage
Source: Orbis database, own calculations.
Notes: The left bar in each year shows the share of MNEs covered by the GMT out of all MNEs. The right bar in each year shows the share of pre-tax profits earned by MNEs covered by GMT out of pre-tax profits by all MNEs. The MNEs covered by GMT are those whose annual revenues exceed $750$ million EUR in at least two of the last four years out of all MNEs. See Appendix B for details.
Against this background, this paper analyzes under which conditions tax competition leads countries to differentiate their tax rates for large and small MNEs in response to the introduction of a GMT. We also study the welfare implications of such split tax system. Finally, we explore whether gradual reforms to the GMT—either through an increase in the GMT rate or an expansion of its coverage—enhance welfare for each country and for the world as a whole.
@@
To answer these questions, we build a simple model where heterogeneous MNEs headquartered in a non-haven country shift profits from there to a set of symmetric tax-haven countries. Countries maximize their welfare, a weighted sum of tax revenues and (for the non-haven) the post-tax income of MNEs’ owners, and they compete for the MNEs’ profits under the GMT with incomplete coverage. Both the non-haven and the havens can decide to commit, in the first stage, to set a uniform tax rate for all their multinationals, or not. In the second stage, all countries choose their tax rates, either uniform or split, where the tax rate applicable for large MNEs must be at or above the GMT level. In the third stage, multinationals make their profit-shifting choices.
@@
Our analysis shows that the introduction of a partial GMT leads to a sequence of tax competition regimes between the non-haven and the havens. If the GMT rate is binding for the tax havens, but still low enough, each of the tax havens will choose to commit to the GMT rate for all MNEs, in order to increase the (uniform) tax rate in the non-haven country. As the GMT rate is further increased, however, the non-cooperative tax equilibrium regime transitions to one where the havens split their tax rates by setting a GMT for large MNEs and a rate lower than the GMT for small MNEs. For sufficiently high GMT rates, a tax regime eventually emerges where the non-haven splits its tax rate as well. This is because a higher GMT rate limits competition for profits from large MNEs, leading the countries to focus instead on profits from small MNEs.
The implications for welfare and tax revenues are as follows. At the regime-switching GMT rates, if one country starts to split its tax rate, the other country discretely loses welfare and tax revenues. In general, however, the non-haven always gains from gradual increases in the GMT rate, whereas the havens weakly gain only for sufficiently low levels of the GMT, and lose otherwise. Gradual increases in the coverage rate of the GMT have similar effects, with the important difference that they never reduce tax revenues in the non-haven country or globally. These results suggest that, although both countries benefit from the introduction of a moderate GMT rate, conflicts of interest will arise once the GMT rate or coverage is increased.
In the final step, we calibrate our model using estimates of profit shifting from a global sample of countries (Tørsløv et al.,, 2023). In our benchmark calibration, the introduction of a $15\%$ GMT rate causes a regime shift: tax haven countries adopt split tax rates, while the non-haven maintains a uniform tax rate above the GMT. This result aligns with the actual responses observed in several tax haven countries, as described above. Regarding welfare and tax revenues, the introduction of the GMT benefits the non-haven while harming havens, leading to net welfare gains worldwide. This conflict of interest between the two groups of countries persists in the two gradual reforms likely to follow: increasing the GMT rate from $15\%$ to $18\%$ and expanding GMT coverage to all MNEs. However, the conflict is less pronounced for the latter reform.
Our model contributes to the recent theoretical literature on the effects of tax competition in the presence of a GMT. In a model with many non-haven and haven countries, Johannesen, (2022) shows that introducing a GMT will always benefit the tax havens, because the GMT reduces the competition between them. In contrast, the welfare effects of the GMT are generally ambiguous for the non-haven country, because of the offsetting changes in tax revenues and firms’ post-tax profits. Hebous and Keen, (2023) also analyze the effects on welfare in a model where two countries differ in terms of both their size and their valuation of public spending. Their quantitative results indicate that GMT levels up to a range of 17 to 20% can constitute strict Pareto improvements, increasing welfare in both high-tax and low-tax countries. Hindriks and Nishimura, (2022) consider governments who choose not only tax rates but also tax enforcement levels. They show that the introduction of the GMT hinders tax enforcement cooperation between the high-tax and the low-tax country, making the high-tax country worse off. Finally, Janeba and Schjelderup, (2023) study a three-country model where two non-haven countries compete for real foreign direct investment (FDI) while simultaneously competing for profit shifting with a third, tax-haven country. In this setting the introduction of the GMT intensifies tax competition for real FDI. While tax revenues in the non-haven countries will still increase if countries compete via corporate tax rates, net revenue gains are zero if competition is via lump-sum subsidies. [^7] Using an international trade model of firms without profit-shifting motives, Nocco et al., (2024) examine a rich set of corporate taxes and subsidies on sales, production, and profits. Specifically, they identify the globally welfare-maximizing combination of non-discriminatory taxes and subsidies including a global minimum tax on firms’ profits.
All these models have in common that the GMT must be levied uniformly on all firms, thus ignoring the existence of a threshold below which the GMT does not apply. In contrast, our model highlights the critical role of this threshold, as it allows countries to use different tax rates on firms with revenues above and below the threshold. This feature links our model to the literature on discriminatory tax competition, which examines whether tax revenues are higher in the non-cooperative tax equilibrium when countries may, or may not, set different tax rates on tax bases with different degrees of international mobility (Janeba and Peters,, 1999; Keen,, 2001; Janeba and Smart,, 2003). [^8] This literature has been extended to cover differences in country size (Bucovetsky and Haufler,, 2007), and imperfect competition in product markets (Gaigné and Wooton,, 2011). A related type of discrimination arises when profit shifting is monitored by governments in a deliberately loose way (Peralta et al.,, 2006). In our setting, tax discrimination will not be profitable for either country in the absence of the GMT, but it will arise in equilibrium if a sufficiently high level of the GMT is imposed. A distinct, but closely related literature has addressed the issue of whether the existence of tax havens—by allowing non-haven countries to tax-discriminate in favor of mobile, multinational firms—raises or reduces tax revenues and welfare in the non-haven countries (Slemrod and Wilson,, 2009; Hong and Smart,, 2010; Johannesen,, 2010; Elsayyad and Konrad,, 2012).
Our analysis proceeds as follows. In Section 2 we set up our model of tax competition and characterize the sub-game perfect Nash equilibria that result in different tax regimes. Section 3 then turns to the welfare effects of introducing a GMT, and of increasing its tax rate and coverage. Section 4 calibrates our model to real-world data and quantifies its effects. Section 5 concludes.
@@ 2.1 Setup
We consider a tax competition model between a non-haven country, indexed by $n$, and a finite set of $H$ symmetric tax-haven countries, indexed by $h$. Tax competition arises between each of the tax havens and the non-haven. There are a large number of heterogeneous, multinational enterprises (MNEs) owned by individuals in the non-haven, each having a headquarters in the non-haven and an affiliate in each of the tax havens. The headquarters in the non-haven makes exogenous profits in this country, $\pi$, which are distributed in the range $[\underline{\pi},\infty)$, with a cumulative distribution function $F(\pi)$. Hence, all real activity occurs in the non-haven, and this is where the “true” profits accrue in its entirety. MNEs can shift a share $\theta\in[0,1]$ of their non-haven headquarters’ profits to the set of tax haven affiliates, in order to maximize their post-tax profits. [^9] This is the simplest possible setting for profit shifting. See e.g., Johannesen, (2022) or Krautheim and Schmidt-Eisenlohr, (2011) for a similar assumption. In particular, this setting avoids the possible non-existence of Nash equilibria that arise in models where the location of heterogeneous firms is endogenous (Baldwin and Okubo,, 2009; Davies and Eckel,, 2010; Haufler and Stähler,, 2013).
We assume that the non-haven and all tax havens have signed the GMT agreement, and are therefore bound to set tax rates of at least the GMT rate, $t_{M}$. [^10] Devereux, (2023) discusses how the specific institutional setting of the GMT, in particular the Under-Taxed Payments Rule (UTPR), gives incentives to both non-haven countries and haven countries to join the GMT agreement, once this has been initiated by a critical number of large non-havens. Note also that we assume the tax base of the GMT to be the same as that of national corporate taxes, thus ignoring extra tax deductions (labelled “substance-based income exclusion”) under the GMT. See Schjelderup and Stähler, (2024) for a study focusing on the latter issue. However, the GMT applies only to MNEs that make exogenous profits no smaller than $\pi_{M}$, called “large MNEs” in our analysis. Therefore, countries remain free to set a lower tax rate than the GMT for all MNEs that make profits less than $\pi_{M}$, called “small MNEs.”
@@
The non-haven country maximizes a welfare function that consists of a weighted sum of tax revenues and the post-tax profits of the heterogeneous multinationals, which are all owned by residents of the non-haven country. Since tax havens do not have any ownership in the firms’ profits, the welfare of tax havens corresponds to their tax revenues.
We consider a three-stage game. In the first stage, the non-haven and haven countries simultaneously choose whether to commit to a single tax rate for all MNEs, or not. [^11] This is related to the literature on endogenous timing, or leadership, in tax competition, where countries commit in a pre-play stage to set taxes early or late. See Kempf and Rota-Graziosi, (2010). To comply with the GMT agreement, this single tax rate has to be equal to, or higher than, the GMT rate. In our main model, we assume that the set of tax havens collectively decides on whether to set a single GMT rate, or split their tax rates. This assumption simplifies the algebra, but it does not change the qualitative nature of our results. [^12] We discuss the alternative of a decentralized first stage decision of the tax havens in Appendix A.2. In the second stage, all countries choose their tax rate(s) in a non-cooperative way. Countries that have committed to a single tax rate in the first stage set a single tax rate, whereas countries that have not committed choose different tax rates for large and small MNEs (where the tax rate on large MNEs has to be at least $t_{M}$). In the final stage, MNEs engage in profit shifting, given the tax rates in the non-haven and the symmetric tax havens. We solve this three-stage game by backward induction. The equilibrium concept we rely on is the sub-game perfect Nash equilibrium, which we call the tax-competition equilibrium.

### v6

@@
A GLOBAL MINIMUM TAX FOR LARGE FIRMS ONLY:
IMPLICATIONS FOR TAX COMPETITION***Andreas Haufler, Seminar for Economic Policy, LMU Munich, Akademiestr. 1, 80799 Munich, Germany. Email: Andreas.Haufler@econ.lmu.de.  Hayato Kato, Graduate School of Economics, the University of Osaka, 1-7 Machikaneyama, Toyonaka, Osaka 560-0043, Japan. Email: hayato.kato@econ.osaka-u.ac.jp.
 Financial support from the German Research Foundation through CRC TRR 190 (Haufler) and the JSPS KEKENHI (JP20H01495, JP22K13388, JP24H00014) (Kato) is gratefully acknowledged. This study is conducted as a part of the Project “Economic Policy Issues in the Global Economy” undertaken at the Research Institute of Economy, Trade and Industry (RIETI). Kato thanks the University of Munich for their hospitality while part of this paper was written. We thank Tomohiro Ara, Real Arai, Ruby Doeleman, Antoine Ferey, Emanuel Hansen, Jim Hines, Niels Johannesen, Makoto Hasegawa, Jota Ishikawa, Mick Keen, Sang-Hyun Kim, Yoshimasa Komoriya, Jonas Löbbing, Mohammed Mardan, Jakob Miethe, Kenichi Nishikata, Hirofumi Okoshi, Pascalis Raimondos, Dirk Schindler, Karl Schulz, Georg Thunecke, and Alfons Weichenrieder for many helpful comments and discussions. Thanks also to seminar and conference participants at the IIPF in Utah State, MPI Berlin, Goettingen, Osaka, Kyoto, Shimane, QUT Brisbane, Ryukoku, Gakushuin, CESifo (Munich & Venice Summer Institute), and RIETI for helpful comments.
Andreas Haufler and Hayato Kato
@@ A GLOBAL MINIMUM TAX FOR LARGE FIRMS ONLY:
The Global Minimum Tax (GMT) is applied only to firms above a certain size threshold, permitting countries to set differential tax rates for small and large firms. We analyse tax competition among multiple tax havens and a non-haven country for heterogeneous multinationals to evaluate the effects of this partial coverage of GMT. Upon the introduction of a moderately low GMT rate, the havens commit to the single uniform GMT rate for all multinationals. However, gradual increases in the GMT rate induce the havens, and subsequently the non-haven, to adopt discriminatory, lower tax rates for small multinationals. Our calibration exercise shows that introducing a GMT rate of 15% results in a regime where only the havens adopt split tax rates. Welfare and tax revenues fall in the havens but rise in the non-haven, yielding a positive net gain worldwide.
Profit shifting by multinational enterprises (MNEs) remains a central challenge for international corporate taxation. Aggregate estimates suggest that more than one-third of foreign-earned corporate profits are shifted to tax havens (Tørsløv et al., , 2023), a pattern also documented in microdata for France (Davies et al., , 2018), the UK (Bilicka, , 2019), and the US (Guvenen et al., , 2022) In response to the resulting revenue losses, the OECD launched its Base Erosion and Profit Shifting (BEPS) initiative (OECD, , 2013), whose core outcome is the introduction of a global minimum tax (GMT), also known as Pillar Two. [^1] See Fuest and Neumeier, (2023); Hebous et al., (2024) for the institutional background. Another outcome of the BEPS initiative is Pillar One, which reallocates taxing rights from the country of a MNE’s residence to the market jurisdiction (Mukunoki et al., , 2025; Ogawa and Tsuchiya, , 2025; Richter, , 2025). In 2021, more than 130 countries including many tax havens agreed on a 15% GMT; by August 2025, 65 countries had drafted or enacted legislation implementing the GMT. [^2] See Tax Foundation, ‘The Latest on the Global Tax Agreement’: https://taxfoundation.org/blog/global-tax-agreement/ (accessed October 1, 2025). Despite announcing after the June 2025 G7 summit that it would not fully comply with the GMT, the US had already implemented a similar tax on Global Intangible Low-Tax Income under the 2017 Tax Cuts and Jobs Act (Chodorow-Reich et al., , 2024).
@@
A limitation of the GMT is that it applies only to ‘large’ multinationals, defined as groups with at least 750 million EUR in annual revenues in two of the last four years. [^3] Since 2016, such groups have been required to file country-by-country reports on their activities, profits, and taxes in each country of operation (Doeleman et al., , 2024, Section 2). The revenue threshold reflects the balance between limiting profit shifting and minimising compliance costs (OECD, , 2022, p.14). According to the Orbis database of Bureau van Dijk, about 30% of MNEs, accounting for roughly 90% of total profits, exceeded this threshold in 2018–2021 and were thus covered by the GMT. [^4] See Supplemental Appendix A. This figure aligns with the OECD’s estimate of 90% profit coverage (OECD, , 2020, # 505, p. 233). Nevertheless, over 300 billion EUR of MNEs’ profits remain outside its scope. Moreover, Orbis tends to oversample large firms (Bajgar et al., , 2020), so the 90% coverage rate should be viewed as an upper bound.
Does this incomplete coverage affect the effectiveness of the GMT? It does not if low-tax countries apply the GMT rate to all firms regardless of size. It does matter, however, if they differentiate their tax rates, applying the GMT rate to firms above the threshold and a lower rate to those below. Such a split is inherent in the regulation: tax havens may implement the GMT by raising their (low) base tax rates only for MNEs above the threshold through a top-up tax. [^5] This is known as the Qualified Domestic Minimum Top-Up Tax (QDMTT) (Devereux, , 2023). This system has been adopted in Bahrain, Ireland, Switzerland and the UAE, among other countries with base statutory corporate tax rates below 15% (PwC, , 2025).
Against this background, we ask when tax competition leads countries to differentiate tax rates for large and small MNEs under the GMT. Does a 15% GMT rate induce such a split, and how large are the resulting revenue and welfare gains? How do these outcomes change with broader GMT coverage?
@@
To answer these questions, we build a simple model where heterogeneous MNEs headquartered in a non-haven country shift profits to tax-haven countries. Countries maximise welfare—a weighted sum of tax revenues and, for the non-haven, the post-tax income of MNEs’ owners—and compete for the MNEs’ profits. In the first stage, both the non-haven and the havens decide whether to commit to a uniform tax rates for all MNEs. In the second stage, countries set tax rates, either uniform or split, with large MNEs taxed at least at the GMT rate. The final stage involves MNEs’ profit-shifting decisions.
@@
Our analysis shows that introducing the GMT with partial coverage generates a sequence of non-cooperative tax equilibrium regimes between the non-haven and the havens. When the GMT rate is binding but low, the havens commit to it for all MNEs to raise the non-haven’s uniform tax rate. As the GMT rate rises, the havens begin to split their rates, applying the GMT to large MNEs and a lower rate to small MNEs. For sufficiently high GMT rates, the non-haven also splits its rate, as limited competition for large MNEs’ profits shifts the focus towards small MNEs.
We calibrate our model using estimates of profit shifting and examine the quantitative effects of GMT reforms. A 15% GMT rate induces the havens to split their tax rates, which is consistent with observed responses in some tax havens including Ireland and Switzerland. It benefits the non-haven while harms the havens, generating net global revenue (welfare) gains of 136 billion USD (83 bUSD), comparable to existing estimates (Baraké et al., , 2022; Hugger et al., , 2024). If the GMT covers all firms, the 15% rate does not produce a split tax system but increases revenues and welfare in the non-haven and globally even more. This global revenue gain would have to be set against the additional compliance costs for multinationals, however.
We contribute to the recent theoretical studies on tax competition under a GMT. They ask how introducing a GMT affects national and global revenues and welfare in the presence of country heterogeneity (Johannesen, , 2022; Hebous and Keen, , 2023); real economic activities (Janeba and Schjelderup, , 2023; Schjelderup and Stähler, , 2024); and endogenous tax-enforcement levels (Hindriks and Nishimura, , 2025). All these models assume the GMT applies uniformly, ignoring the firm-size threshold below which firms are exempt from the GMT. Our model highlights this threshold, allowing countries to tax firms above and below it differently. [^6] This links our model to the literature on discriminatory tax competition, where countries can tax firms differently by their mobility (Janeba and Peters, , 1999; Keen, , 2001; Bucovetsky and Haufler, , 2007; Gaigné and Wooton, , 2011).
Our simple quantitative exercise complements recent evaluations of corporate taxation using trade and multinational models (Wang, , 2020; Ferrari et al., , 2023; Shen, , 2024; Dyrda et al., , 2024), where taxes are assumed to be exogenous and/or uniform across firms. By contrast, we abstract from firms’ real activities and focus on the strategic role of size-dependent taxes.
Two empirical papers (Bachas et al., , 2025; Clifford et al., , 2025) closely related to ours examine which firms fall within the scope of the GMT using firm/affiliate level data for 16 countries and Germany, respectively. Their findings that GMT revenue gains depend strongly on its coverage emphasizes the relevance of our analysis.
@@ 1.1 Setup
We consider tax competition between a non-haven country, indexed by $n$, and a finite set of $H$ symmetric tax-haven countries, indexed by $h$. There is a continuum of heterogeneous MNEs which are headquartered in the non-haven and have affiliates in all havens. Headquarters earn (‘true’) exogenous profits $\pi\in[\underline{\pi},\infty)$ with distribution $F(\pi)$, arising from real activity in the non-haven. [^7] Firm heterogeneity could result from (executives’) productivity (e.g., Melitz, , 2003; Lippi and Schivardi, , 2014). MNEs can shift some of these profits to the havens to maximise their post-tax profits. In the following, we assume collective decision-making by the havens. This assumption is relaxed in the quantitative analysis later (see also footnote 16) and further examined in Supplemental Appendix B.
We assume the non-haven and all havens have signed the GMT agreement and must set tax rates no lower than the GMT rate $t_{M}$. [^8] Devereux, (2023) analyses how the GMT’s institutional design, especially the Under-Taxed Payments Rule (UTPR), incentivizes non-haven and haven countries to join once a critical mass of large non-havens participates. The GMT applies only to MNEs with profits $\pi\geq\pi_{M}$ (‘large MNEs’), allowing countries to set lower rates for firms with $\pi<\pi_{M}$ (‘small MNEs’).
@@
The non-haven maximises a welfare function that consists of a weighted sum of tax revenues and the post-tax profits of MNEs, which are all owned by residents of the non-haven. Since havens do not have any ownership in the firms’ profits, the welfare of tax havens corresponds to their tax revenues.
We analyse a three-stage game. In the first stage, the non-haven and havens simultaneously decide whether to commit to a single tax rate for all MNEs—at or above the GMT rate—or to allow differentiated rates. [^9] The ability to commit, a realistic feature of international tax relations, is standard both in the studies cited in footnote 6 and in related work on endogenous timing in games (Huck et al., , 2002; Kempf and Rota-Graziosi, , 2010). The choice between a uniform and a split tax system is more fundamental, and less frequently revised, than the choice of tax rates. The havens decide collectively on whether to unify or split their rates. In the second stage, countries set their tax rates non-cooperatively. Those that committed choose a single rate; others set different rates for large and small MNEs. In the final stage, MNEs shift profits given these tax rates. The game is solved by backward induction, yielding a sub-game perfect Nash equilibrium, referred to as the tax-competition equilibrium.